# Structure and dynamics of electron-phonon coupled systems using neural quantum states

### v1

@@ 1 Introduction
Since Landau proposed the concept of the self-trapping of an electron via lattice distortions, there has been extensive theoretical investigation into eph interactions.Alexandrov and Devreese (2010); Franchini et al. (2021); Giustino (2017) This exploration has entailed the development of simple lattice and continuum models, such as those pioneered by HolsteinHolstein (1959, 1959) and Fröhlich,Fröhlich (1954) alongside methods to solve the resulting eph problem. Various strategies for tackling these models include path integralFeynman (1955); Kornilovitch (1998); Beyl et al. (2018); Lee et al. (2021); Cohen-Stead et al. (2022) and diagrammatic methods,Prokof’ev and Svistunov (1998); Berciu (2006) variationalLee et al. (1953); Toyozawa (1961); Brown et al. (1997); Alder et al. (1997); Ohgoe and Imada (2014); Hohenadler et al. (2004); Karakuzu et al. (2017); Ferrari et al. (2020) and perturbative techniques,Marsiglio (1995); Hohenadler et al. (2003) density matrix renormalization group (DMRG) approachesJeckelmann and White (1998), as well as various flavors of exact diagonalizationBonča et al. (1999); Wang et al. (2020). Although some of these approaches are numerically exact in some domains, their applicability across different types of Hamiltonians and parameter ranges is constrained in practice due to potentially high computational costs. Recently, significant progress has been made in constructing and studying ab initio models based on density functional perturbation theory (DFPT).Baroni et al. (2001); Sjakste et al. (2015); Giustino (2017); Poncé et al. (2016); Zhou et al. (2021); Verdi and Giustino (2015) Methods initially developed for the investigation of lattice models are now being adapted to this ab initio setting.Lee et al. (2021); Lafuente-Bartolome et al. (2022) Systematically improvable calculations towards the exact limit have not yet been reported. In addition to allowing quantitative comparison with experiments and predictive power, such calculations also help assess the accuracy of approximate but more affordable computational methods.
@@
The flexibility and accuracy of NQS come at the expense of having to perform stochastic nonlinear optimization, presenting challenges more severe than those encountered in deterministic methods.Cai and Liu (2018); Westerhout et al. (2020); Szabó and Castelnovo (2020) Good results require careful development of heuristics for the choice of representations and learning strategies, which have to be tailored to the problem domain. In this work, we present our efforts to address these challenges for eph systems. We find that the NQS approaches we put forward here are competitive with the best exact approaches for standard polaron problems, while providing facile access to highly accurate spectral properties and the ability to flexibly treat complex problems such as those with correlated electrons and ab initio parametrized Hamiltonians.
In the following sections, we present the details of Hamiltonians, wave functions, and the methods used to calculate their properties (Sec. 2). We then present the results of our calculations on a variety of eph systems to demonstrate the utility of our approach (Sec. 3). Finally, we conclude with a discussion of our results and future directions (Sec. 4).
@@ 2.1 Hamiltonians
Here $c_{n,k}$ are electronic annihilation operators for an electron with crystal momentum $k$ in band $n$, and $b_{\nu q}$ are phonon annihilation operators with crystal momentum $q$ in band $\nu$. $\epsilon_{k}$ are electronic energies, $V_{k_{1}k_{2}k_{3}}^{ijmn}$ are electron-electron interactions, $\omega_{\nu,q}$ are phonon energies, and $g_{mn\nu kq}$ are eph couplings. We have omitted spin indices for brevity. In an ab initio setting, the band energies and interaction terms can be obtained from DFPT calculations.
@@ 2.2 Wave functions
$\psi(\textbf{n})=\frac{1}{N}\sum_{g}c_{g}\psi(g\textbf{n}),$ (7)
@@
where $P_{k}$ denotes a projector onto the subspace of fixed momentum $k$, which acts as in Eq. 7. Translationally symmetrized NN states can thus be thought of as a generalization of the Toyozawa ansatz. Note that this wave function can be evaluated and optimized deterministically at a polynomial cost, but we use it within VMC here. The low computational cost of its evaluation allows us to assess finite-size effects in cases where calculations with NN states are expensive on large systems.
@@ 2.4 Real-frequency Green’s functions and excited states
While it is possible to sample the amplitude square of the ground state to obtain $\mathbf{H}$ and $\mathbf{S}$, this can lead to noisy estimates, as $|\mu\rangle$ states may have support on configurations with vanishing contributions to $|\psi_{0}\rangle$. Moreover, noise in the unbiased estimates of these matrix elements leads to a bias in the estimates of the energy spectrum. We instead employ an alternative approach, termed reweighting in Ref.  87, which uses $\sum_{\nu}|\langle n|\psi_{\mu}\rangle|^{2}$ as the sampling function instead. The two approaches lead to identical results in the limit of infinite sampling, but we found the reweighting method to perform significantly better for a fixed sampling effort. Details and comparisons of these sampling approaches are provided in Appendix B. Due to stochastic sampling, the estimated metric $S$ is not necessarily positive definite and needs to be regularized. We diagonalize the sampled metric and throw away states with eigenvalues below a small threshold.
@@
Figure 1: Convergence of error of the NQS ground state energy for the Holstein (left) and Bond (right) model polarons on a 30 site chain with respect to the number of hidden neurons ($n_{H}$) with $\omega_{0}=0.5$.
This approach of using derivative states has been used with many classic wave function theories, often termed as linear response or equation of motion (EOM) methods.McWeeny (1989) The Tamm-Dancoff approximation, which is formulated in the tangent space of the Hartree-Fock state, is a well-known example. A linear response DMRG theory has been developed in close analogy to Hartree-Fock.Dorando et al. (2009); Wouters et al. (2013); Haegeman et al. (2013) Similarly to NQS, linear response DMRG can also be systematically improved by increasing the bond dimension of the matrix product state. In VMC, an EOM theory based on geminal wave functions was presented in Ref.  91. A few studies have obtained dynamical information through the use of a basis constructed from physically relevant excitations on top of the variational ground state.Ido et al. (2020); Ferrari et al. (2020) Derivative basis states have the advantage of being computationally cheaper for the calculation of the required matrix elements. For example, consider the following element used in the sampling of the Hamiltonian matrix
@@ 3 Results
We present a numerical analysis of the performance of NQS in polaron, bipolaron, and many-electron systems. First, we consider the ability of these states to represent ground state structure for different kinds of eph coupling. We also calculate the binding energy of the hole polaron in lithium fluoride (LiF) from first principles. In the second part, we assess the accuracy of our LR-VMC approach based on NQS to capture spectral properties of eph systems. We use three-layer MLPs (input, hidden, output) for the radial and phase part of the NQS in all cases, unless stated otherwise. The code used to perform the VMC calculations is available in a public repository.vmc
@@ 3.1 Ground state properties
We start with the question of how many hidden neurons are required to accurately represent the ground state of Holstein and Bond model polarons at different coupling strengths. For the Holstein model, the phase function was set to one, whereas for the Bond model, we use MLPs with identical structures for both the radial and phase parts of the wave function in momentum space. Fig. 1 shows the errors in the ground state energies for a 30 site chain with $\omega=0.5$. We use energies converged with respect to the number of hidden neurons as a reference. We have verified that these values agree very well with DMRG in the case of the Holstein model and numerically exact energies reported in literatureCarbone et al. (2021) for the Bond model wherever available. We ensured that the ground state energy in all cases is converged with respect to the number of lattice sites. For the Holstein model, the error is largest for a given number of neurons at intermediate coupling around the self-trapping crossover. This is intuitively sensible since simple weak and strong coupling ansatzes describe the regions away from the crossover point very well. The ansatz with a single hidden neuron, which is equivalent to the Toyozawa wave function, performs very well for the Holstein model. Convergence with respect to the number of hidden neurons is achieved very quickly, with the energies close to exact with just 10 hidden neurons. For the Bond model, on the other hand, the error increases with coupling for a fixed number of hidden neurons. The errors for the same number of hidden neurons are much larger compared to the Holstein model. This model does not exhibit a self-trapping crossover, and the number of hidden neurons required to obtain the same error in ground state energy increases with the size of the coupling. This also coincides with there not being a simple strong coupling ansatz that describes the strong coupling limit of the Bond model. We note that these observations also hold for calculations in the site basis and in higher dimensions. The number of hidden neurons for a given energy accuracy does not scale with the lattice size.
@@
Figure 3: Ground state energy of the Holstein, Bond, and SSH polarons on a two-dimensional lattice for $\omega_{0}=0.5$ (upper curves) and $\omega_{0}=3$ (lower curves). DQMC energies taken from Ref.  95
To demonstrate the robustness of our optimization protocol, we present a calculation on the ground state band of a modified Holstein model polaron with dispersive phonons. This system was studied using variational exact diagonalization (VED) in Ref.  96. We use the phonon dispersion given by
@@
Fig. 3 illustrates the behavior of the polaron ground state energy on two-dimensional lattices. We used a $10\times 10$ lattice for all calculations, which was confirmed to be sufficiently large to reach convergence within stochastic error bars for all cases presented here. For the Holstein model, similar to the one-dimensional chain considered previously, we find that the ground state energy is readily converged with a small number of hidden neurons. In fact, the Toyozawa wave function (equivalent to one hidden neuron) is very accurate in this case even on the two-dimensional lattice. We see the self-trapping crossover in the Holstein model clearly in this plot. The Bond and SSH models are considerably more difficult to solve for the NQS approach indicating the complexity of the off-diagonal coupling. We compare our energies to DQMC results reported in Ref.  95. For the Bond model, we are able to converge NQS energies to the DQMC results with a larger number of hidden neurons, around 100 for the largest couplings. The Toyozawa ansatz exhibits substantial errors that increase with the size of the coupling. The SSH model at large couplings has its ground state at nonzero $k$ due to the negative next-nearest neighbor hopping amplitude induced by eph coupling. This behavior has been attributed to the unphysical linear nature of the interaction at strong couplings and was shown to disappear with more realistic nonlinear couplings.Prokof’ev and Svistunov (2022) Nevertheless, we find that the NQS is able to capture this shift of the ground state off the band center accurately. In this regime of stronger coupling, optimization becomes very slow and the ansatz requires a large number of hidden neurons to reach convergence. In these cases, we extrapolate our energies with respect to the number of hidden neurons (see Appendix A). We do not show the Toyozawa ansatz energies in this case because we were unable to converge them reliably.
@@
Figure 5: Binding energy of the LiF hole polaron obtained using NQS. “LO only” indicates calculations performed with only the longitudinal optical phonon mode, whereas the “All modes” energies are obtained with all three acoustic and optical modes. The star indicates the binding energy extrapolated to the thermodynamic limit.
@@
Figure 6: Binding energy of the Bond and SSH bipolarons on a two-dimensional square lattice with $\omega_{0}=0.5$ and $\lambda=0.5$ (lower curve) and 1 (upper curve) as a function of on-site repulsion $U$. DQMC results for the Bond model taken from Ref.  97.
@@
For the final polaron example, we calculate the hole polaron binding energy in lithium fluoride (LiF) from first principles.Sio et al. (2019); Lee et al. (2021) This is a polar ionic crystal with a large band gap (14.2 eV experimental optical gap, 8.9 eV calculated PBE KS gap).Sio et al. (2019) We obtain the Hamiltonian parameters from DFT calculations using Quantum EspressoGiannozzi et al. (2017) with the EPW package.Poncé et al. (2016) We include three valence bands arising mainly from the $p$ orbitals of fluorine in our calculation. These are coupled to three optical and three acoustic mode vibrations. Due to the polar nature of this system, the hole coupling to the longitudinal optical (LO) mode is the strongest and largely of the Fröhlich type. Due to the very strong coupling of the hole with phonons, we expect the ground state to be readily describable with a strong coupling ansatz. This is borne out in our calculations shown in Fig. 5. We calculate ground state energies on progressively larger $k$ point grids ranging from $3\times 3\times 3$ to $13\times 13\times 13$. Due to the large number of input sites in the NQS state, we are restricted in the number of hidden neurons we can employ in this calculation. Therefore as a validation of our ansatz, we also calculate energies on a smaller model including just the LO phonon mode in the Hamiltonian, which contributes the bulk of the binding energy. We find that increasing the number of hidden neurons provides modest improvements (of the order of only 10 meV) over the Toyozawa ansatz as expected in a strong density coupling case. Using a crude two-point extrapolation we obtain a binding energy of 1.97 eV in the thermodynamic limit. The largest uncertainty in this number, which is not reported, arises simply due to the crude nature of the extrapolation from small grid sizes. A future study will be devoted to a more careful extrapolation of this value. This binding energy is in good agreement with strong coupling perturbation theory calculations and results obtained from a novel all-coupling wave function ansatz,Robinson et al. as well as values reported in the literature using Landau-Pekar theory.Sio et al. (2019) With improvements in our numerical methodology, including the use of locality of interactions and low-rank compression,Luo et al. (2024) we expect to be able to perform calculations on even larger grids in the future.

### v2

@@ 1 Introduction
Since Landau proposed the concept of the self-trapping of an electron via lattice distortions, there has been extensive theoretical investigation into eph interactions.Alexandrov and Devreese (2010); Franchini et al. (2021); Giustino (2017) This exploration has entailed the development of simple lattice and continuum models, such as those pioneered by HolsteinHolstein (1959, 1959) and Fröhlich,Fröhlich (1954) alongside methods to solve the resulting eph problem. Various strategies for tackling these models include path integralFeynman (1955); Kornilovitch (1998); Beyl et al. (2018); Lee et al. (2021); Cohen-Stead et al. (2022); Malkaruge Costa et al. (2023); Li and Johnston (2023) and diagrammatic methods,Prokof’ev and Svistunov (1998); Berciu (2006) variationalLee et al. (1953); Toyozawa (1961); Brown et al. (1997); Alder et al. (1997); Ohgoe and Imada (2014); Hohenadler et al. (2004); Karakuzu et al. (2017); Ferrari et al. (2020) and perturbative techniques,Marsiglio (1995); Hohenadler et al. (2003) density matrix renormalization group (DMRG) approachesJeckelmann and White (1998), as well as various flavors of exact diagonalizationBonča et al. (1999); Wang et al. (2020). Although some of these approaches are numerically exact in some domains, their applicability across different types of Hamiltonians and parameter ranges is constrained in practice due to potentially high computational costs. Recently, significant progress has been made in constructing and studying ab initio models based on density functional perturbation theory (DFPT).Baroni et al. (2001); Sjakste et al. (2015); Giustino (2017); Poncé et al. (2016); Zhou et al. (2021); Verdi and Giustino (2015) Methods initially developed for the investigation of lattice models are now being adapted to this ab initio setting.Lee et al. (2021); Lafuente-Bartolome et al. (2022) Systematically improvable calculations towards the exact limit have not yet been reported. In addition to allowing quantitative comparison with experiments and predictive power, such calculations also help assess the accuracy of approximate but more affordable computational methods.
Figure 1: Schematic illustration of the neural quantum state (NQS) approach for electron-phonon coupled systems showing steps in the VMC sampling, optimization, and calculations of dynamical properties.
@@
The flexibility and accuracy of NQS come at the expense of having to perform stochastic nonlinear optimization, presenting challenges more severe than those encountered in deterministic methods.Cai and Liu (2018); Westerhout et al. (2020); Szabó and Castelnovo (2020) Good results require careful development of heuristics for the choice of representations and learning strategies, which have to be tailored to the problem domain. In this work, we present our efforts to address these challenges for eph systems. We find that the NQS approaches we put forward here are competitive with the best exact approaches for standard polaron problems, while providing facile access to highly accurate spectral properties and the ability to flexibly treat complex problems such as those with correlated electrons and ab initio parametrized Hamiltonians. In the following sections, we present the details of Hamiltonians, wave functions, and the methods used to calculate their properties (Sec. 2). We then present the results of our calculations on a variety of eph systems to demonstrate the utility of our approach (Sec. 3). Finally, we conclude with a discussion of our results and future directions (Sec. 4).
@@ 2.1 Hamiltonians
Here $c_{n,k}$ are electronic annihilation operators for an electron with crystal momentum $k$ in band $n$, and $b_{\nu q}$ are phonon annihilation operators with crystal momentum $q$ in band $\nu$. $\epsilon_{k}$ are electronic energies, $V_{k_{1}k_{2}k_{3}}^{ijmn}$ are electron-electron interactions, $\omega_{\nu,q}$ are phonon energies, and $g_{mn\nu kq}$ are eph couplings. We have omitted spin indices for brevity. In an ab initio setting, we obtain the band energies and interaction terms from DFPT calculations.Giustino (2017); Poncé et al. (2016); Zhou et al. (2021)
@@ 2.2 Wave functions
$\psi^{S}(\textbf{n})=\frac{1}{N}\sum_{g}c_{g}\psi(g\textbf{n}),$ (7)
@@
where $P_{k}$ denotes a projector onto the subspace of fixed momentum $k$, which acts as in Eq. 7. Translationally symmetrized NN states can thus be thought of as a generalization of the Toyozawa ansatz. Note that this wave function can be evaluated and optimized deterministically at a polynomial cost, but we use it within VMC here. The low computational cost of its evaluation allows us to assess finite-size effects in cases where calculations with NN states are expensive on large systems. Other generalizations of the Davydov ansatz like the global-localBrown et al. (1997) and delocalized D1 ansatzesSun et al. (2013) yield more accurate results compared to Toyozawa, especially for non-local couplings.Lee It would be interesting to explore the possibility of basing NN states on these wave functions in future work.
@@ 2.4 Real-frequency Green’s functions and excited states
While it is possible to sample the amplitude square of the ground state to obtain $\mathbf{H}$ and $\mathbf{S}$, this can lead to noisy estimates, as $|\mu\rangle$ states may have support on configurations with vanishing contributions to $|\psi_{0}\rangle$. Moreover, noise in the unbiased estimates of these matrix elements leads to a bias in the estimates of the energy spectrum. We instead employ an alternative approach, termed reweighting in Ref. 91, which uses $\sum_{\nu}|\langle n|\psi_{\mu}\rangle|^{2}$ as the sampling function instead. The two approaches lead to identical results in the limit of infinite sampling, but we found the reweighting method to perform significantly better for a fixed sampling effort. Details and comparisons of these sampling approaches are provided in Appendix B. Due to stochastic sampling, the estimated metric $S$ is not necessarily positive definite and needs to be regularized. We diagonalize the sampled metric and throw away states with eigenvalues below a small threshold.
@@
Figure 2: Convergence of error of the NQS ground state energy for the Holstein (left) and Bond (right) model polarons on a 30 site chain with respect to the number of hidden neurons ($n_{H}$) with $\omega_{0}=0.5$. We use DMRG energies converged with respect to the number of phonons as a reference.
This approach of using derivative states has been used with many classic wave function theories, often termed as linear response or equation of motion (EOM) methods.McWeeny (1989) The Tamm-Dancoff approximation, which is formulated in the tangent space of the Hartree-Fock state, is a well-known example. A linear response DMRG theory has been developed in close analogy to Hartree-Fock.Dorando et al. (2009); Wouters et al. (2013); Haegeman et al. (2013) Similarly to NQS, linear response DMRG can also be systematically improved by increasing the bond dimension of the matrix product state. In VMC, an EOM theory based on geminal wave functions was presented in Ref. 95. A few studies have obtained dynamical information through the use of a basis constructed from physically relevant excitations on top of the variational ground state.Ido et al. (2020); Ferrari et al. (2020) Derivative basis states have the advantage of being computationally cheaper for the calculation of the required matrix elements. For example, consider the following element used in the sampling of the Hamiltonian matrix
@@ 3 Results
We present a numerical analysis of the performance of NQS in polaron, bipolaron, and many-electron systems. First, we consider the ability of these states to represent ground state structure for different kinds of eph coupling. We also calculate the binding energy of the hole polaron in lithium fluoride (LiF) from first principles. In the second part, we assess the accuracy of our LR-VMC approach based on NQS to capture spectral properties of eph systems. We use three-layer MLPs (input, hidden, output) for the radial and phase part of the NQS in all cases, unless stated otherwise. The code used to perform the VMC calculations is available in a public repository.vmc DMRG results used for reference were obtained using the ITensor library.Fishman et al. (2022)
@@ 3.1 Ground state properties
3.1.1 Convergence with number of parameters
We start with the question of how many hidden neurons are required to accurately represent the ground state of Holstein and Bond model polarons at different coupling strengths. For the Holstein model, the phase function was set to one, whereas for the Bond model, we use MLPs with identical structures for both the radial and phase parts of the wave function in momentum space. Fig. 2 shows the relative percent errors in the ground state energies for a 30 site chain with $\omega=0.5$. We use DMRG energies converged with respect to the number of phonons as a reference. For the Holstein model, the error is largest for a given number of neurons at intermediate coupling around the self-trapping crossover. This is intuitively sensible since simple weak and strong coupling ansatzes describe the regions away from the crossover point very well. The ansatz with a single hidden neuron, which is equivalent to the Toyozawa wave function, performs very well for the Holstein model. Convergence with respect to the number of hidden neurons is achieved very quickly, with the energies close to exact with just 10 hidden neurons. For the Bond model, on the other hand, the error increases with coupling for a fixed number of hidden neurons. The errors for the same number of hidden neurons are much larger compared to the Holstein model. This model does not exhibit a self-trapping crossover, and the number of hidden neurons required to obtain the same error in ground state energy increases with the size of the coupling. This also coincides with there not being a simple strong coupling ansatz that describes the strong coupling limit of the Bond model. We note that these observations also hold for calculations in the site basis and in higher dimensions. The number of hidden neurons for a given energy accuracy does not scale with the lattice size.
@@ 3.1.1 Convergence with number of parameters
Figure 4: Ground state energy of the Holstein, Bond, and SSH polarons on a two-dimensional lattice for $\omega_{0}=0.5$ (upper curves) and $\omega_{0}=3$ (lower curves). DQMC energies taken from Ref. 99
3.1.2 Dispersive phonons
To demonstrate the robustness of our optimization protocol, we present a calculation on the ground state band of a modified Holstein model polaron with dispersive phonons. This system was studied using variational exact diagonalization (VED) in Ref. 100. We use the phonon dispersion given by
@@ 3.1.2 Dispersive phonons
3.1.3 Two-dimensional polarons
Fig. 4 illustrates the behavior of the polaron ground state energy on two-dimensional lattices. We used a $10\times 10$ lattice for all calculations, which was confirmed to be sufficiently large to reach convergence within stochastic error bars for all cases presented here. For the Holstein model, similar to the one-dimensional chain considered previously, we find that the ground state energy is readily converged with a small number of hidden neurons. In fact, the Toyozawa wave function (equivalent to one hidden neuron) is very accurate in this case even on the two-dimensional lattice. We see the self-trapping crossover in the Holstein model clearly in this plot. The Bond and SSH models are considerably more difficult to solve for the NQS approach indicating the complexity of the off-diagonal coupling. We compare our energies to DQMC results reported in Ref. 99. For the Bond model, we are able to converge NQS energies to the DQMC results with a larger number of hidden neurons, around 100 for the largest couplings. The Toyozawa ansatz exhibits substantial errors that increase with the size of the coupling. The SSH model at large couplings has its ground state at nonzero $k$ due to the negative next-nearest neighbor hopping amplitude induced by eph coupling. This behavior has been attributed to the unphysical linear nature of the interaction at strong couplings and was shown to disappear with more realistic nonlinear couplings.Prokof’ev and Svistunov (2022) Nevertheless, we find that the NQS is able to capture this shift of the ground state off the band center accurately. In this regime of stronger coupling, optimization becomes very slow and the ansatz requires a large number of hidden neurons to reach convergence. In these cases, we extrapolate our energies with respect to the number of hidden neurons (see Appendix A). We do not show the Toyozawa ansatz energies in this case because we were unable to converge them reliably.
@@ 3.1.3 Two-dimensional polarons
Figure 6: Binding energy of the LiF hole polaron obtained using NQS. “LO only” indicates calculations performed with only the longitudinal optical phonon mode, whereas the “All modes” energies are obtained with all three acoustic and optical modes. The star indicates the binding energy extrapolated to the thermodynamic limit using a quadratic fit.
Figure 7: Binding energy of the LiF electron polaron obtained using NQS. The star indicates the binding energy extrapolated to the thermodynamic limit using a quadratic fit. Agreement between the fits using the largest 4 and 5 grids simulated demonstrate the fidelity of this fitting procedure.
@@
Figure 8: Binding energy of the Bond and SSH bipolarons on a two-dimensional square lattice with $\omega_{0}=0.5$ and $\lambda=0.5$ (lower curve) and 1 (upper curve) as a function of on-site repulsion $U$. DQMC results for the Bond model taken from Ref. 101.
3.1.4 Ab initio calculations of polaron binding energies in LiF
For the final polaron example, we calculate the hole and electron polaron binding energies in lithium fluoride (LiF) from first principles.Sio et al. (2019); Lee et al. (2021) This is a polar ionic crystal with a large band gap (14.2 eV experimental optical gap, 8.9 eV calculated PBE KS gap).Sio et al. (2019) We obtain the Hamiltonian parameters from DFT calculations using Quantum EspressoGiannozzi et al. (2017) with the EPW package.Poncé et al. (2016) We include three valence bands arising mainly from the $p$ orbitals of fluorine in our hole polaron calculation, while one conduction band was included in the electron polaron calculation. Higher energy bands may in principle be included, but are omitted here. Previous work has indicated that these bands provide the dominant contribution to the binding energy, Sio et al. (2019); Lee et al. (2021) however inclusion of more remote bands may provide quantitative contributions the polaron energetics. The relevant bands are coupled to three optical and three acoustic mode vibrational phonon branches. Due to the polar nature of this system, the hole coupling to the longitudinal optical (LO) mode is the strongest and largely of the Fröhlich type.
@@ 3.1.4 Ab initio calculations of polaron binding energies in LiF
Due to the very strong coupling of the hole with phonons, we expect the hole polaron ground state to be readily describable with a strong coupling ansatz. This is borne out in our calculations shown in Fig. 6. We calculate ground state energies on progressively larger $k$ point grids ranging from $3\times 3\times 3$ to $13\times 13\times 13$. Due to the large number of input sites in the NQS state, we are restricted in the number of hidden neurons we can employ in this calculation. Therefore as a validation of our ansatz, we also calculate energies on a smaller model including just the LO phonon mode in the Hamiltonian, which contributes the bulk of the binding energy. We find that increasing the number of hidden neurons provides modest improvements (of the order of only 10 meV) over the Toyozawa ansatz as expected in a strong density coupling case. Using a quadratic fit of the energies obtained for the three largest grids we obtain a binding energy of 2.14 eV in the thermodynamic limit. Given the uncertainties in the extrapolation procedures due to computational limitations of the grid sizes employed, it is satisfying that our value of the binding energy is in reasonable agreement with strong coupling perturbation theory calculations (1.96 eV) and results obtained from a novel all-coupling wave function ansatz (1.94 eV),Robinson et al. as well as a calculation using many-body Green’s function theory (2.20 eV).Lafuente-Bartolome et al. (2022)
The electron polaron is weakly coupled and more delocalized, requiring larger $k$-grids to converge the binding energy. We performed NN-VMC calculations on grids of sizes up to $21\times 21\times 21$. Fig. 7 shows the binding energy of the electron polaron as a function of the inverse grid size. We used a quadratic fit to extrapolate to the thermodynamic limit in this case as we found it to model the data better. The quadratic fits to the largest 4 and 5 grids produce similar extrapolated energies within 10 meV of each other. The five-point estimate of 0.39 eV is lower than the 0.6 eV obtained using a Green’s function method in Ref. 53, possibly due to the inclusion of contributions from higher conduction bands and quadratic coupling in Ref. 53, or simply due to the fact that the approach of Ref. 53 is approximate.
The largest uncertainty in these binding energies, which are not reported, arises simply due to the nature of the extrapolation from small grid sizes. For example, in the case of the hole polaron, a linear extrapolation using the two largest grids leads to an extrapolated binding energy of 1.97 eV compared to the 2.14 eV obtained using a quadratic fit. A future study will be devoted to a more careful extrapolation of these values. With improvements in our numerical methodology, including the use of locality of interactions and low-rank compression,Luo et al. (2024) we expect to be able to perform calculations on even larger grids in the future.
3.1.5 Bond and SSH model bipolarons